# SDFLoRA: Selective Dual-Module LoRA for Federated Fine-tuning with Heterogeneous Clients

### v1

@@ Abstract
Federated learning (FL) for large language models (LLMs) has attracted increasing attention for enabling privacy-preserving adaptation across distributed data, where parameter-efficient methods such as LoRA are widely adopted to reduce communication and memory costs. Despite these advances, practical FL deployments exhibit rank heterogeneity, as clients often use different low-rank configurations, making direct aggregation of LoRA updates biased and unstable. Existing solutions enforce unified ranks or align heterogeneous updates into a shared subspace, which over-constrains client-specific semantics, limits personalization, and provides weak protection of local client information under differential privacy noise. To address this gap, we propose Selective Dual-module Federated LoRA (SDFLoRA), which decomposes each client adapter into a global module for transferable knowledge and a local module for client-specific adaptations. The global module is selectively aligned and aggregated across clients, while local modules remain private, enabling robust learning under rank heterogeneity and privacy-aware optimization by injecting differential privacy noise exclusively into the global module. Experiments on GLUE benchmarks show that SDFLoRA outperforms representative federated LoRA baselines and achieves a better utility–privacy trade-off.
@@ 1 Introduction
Parameter-efficient federated fine-tuning has become a practical route for adapting large language models (LLMs) when data are distributed and privacy-sensitive. By optimizing only lightweight adapters, clients avoid full-model communication and memory overhead, enabling deployments in domains such as healthcare, finance, and on-device intelligence. Among PEFT approaches, low-rank adaptation (LoRA) Hu et al. (2022) is the dominant choice in federated LLM adaptation and is widely adopted in recent FL-LLM systems Wang et al. (2024); Yang et al. (2025).
Despite these advances, practical FL deployments could exhibit strong heterogeneity, as clients originate from different domains, tasks, and user populations. Consequently, LoRA updates naturally contain both transferable directions shared across clients and client-specific directions that encode local semantics and personalization. However, existing federated LoRA pipelines typically enforce global alignment and aggregation over all update directions, which suppresses client-specific adaptations and limits personalization. This issue is further exacerbated by rank heterogeneity, where clients adopt different low-rank configurations due to diverse system constraints, making naïve aggregation even more biased and unstable.
Existing solutions mainly follow two directions. One line enforces a unified rank across all clients, which restricts flexibility and wastes resources on constrained devices. Another line aligns heterogeneous LoRA updates into a shared subspace to enable aggregation, but it implicitly aligns all directions, including client-specific components that should remain private and unshared, thereby further degrading personalization. Meanwhile, differential privacy is widely introduced in FL to protect local client information; yet, when structurally mixed updates are indiscriminately aligned, injecting noise into the entire update perturbs client-specific components that should remain private, leading to unnecessary utility loss and a degraded utility–privacy trade-off.
Motivated by these observations, we propose SDFLoRA, which decomposes each client adapter into a global module that captures transferable knowledge and a local module that preserves client-specific adaptations. Only the global module participates in selective subspace alignment and aggregation across clients, while local modules remain private and unaggregated, thereby handling rank heterogeneity without over-constraining personalization. Moreover, this structure enables privacy-aware optimization by injecting noise exclusively into the global module, protecting shared information while avoiding unnecessary perturbations on purely local adaptations, and yielding a better utility–privacy trade-off.
@@
  • We propose the Selective Dual-module Federated LoRA (SDFLoRA) approach for federated LLM fine-tuning, which decomposes each client adapter into a global module for transferable knowledge and a local module for client-specific adaptations, and performs selective subspace aggregation only on the global module to address rank heterogeneity without over-constraining personalization.
  • We develop a privacy-aware optimization scheme in SDFLoRA by injecting differential privacy noise exclusively into the global module while keeping local modules private and unaggregated, improving robustness under DP noise.
  • Experiments on GLUE benchmarks with heterogeneous-rank clients show that SDFLoRA outperforms representative federated LoRA baselines and achieves a better utility–privacy trade-off, highlighting the importance of structural awareness in federated low-rank adaptation.
@@ 2 Related Work
Our work mainly relates to aspects including (i) federated optimization under heterogeneity, (ii) parameter-efficient fine-tuning and federated LoRA, and (iii) personalization and privacy in federated learning.
@@ 2.1 Federated Optimization under Client Heterogeneity
Federated learning (FL) enables collaborative model training without centralizing data by alternating between local optimization and server-side aggregation. Federated Averaging (FedAvg) McMahan et al. (2017) is the canonical approach, but its simple weighted averaging mechanism often leads to unstable convergence under non-IID data distributions, partial client participation, and multiple local update steps, a phenomenon commonly referred to as client drift.
To mitigate heterogeneity-induced optimization challenges, a lot of work has proposed algorithmic modifications from different perspectives. FedProx Li et al. (2020) introduces a proximal regularization term to constrain local updates from deviating excessively from the global model. FedDyn Durmus et al. (2021) and FedDC Gao et al. (2022) further address client drift by incorporating dynamic correction terms that explicitly decouple local objectives from the global optimization trajectory. Another line of work focuses on correcting biases caused by heterogeneous local computation. FedNova Wang et al. (2020) normalizes client updates according to their effective number of local steps, while SCAFFOLD Karimireddy et al. (2020b) reduces update variance through control variates shared between the server and clients. From an optimization-centric viewpoint, MIME and MIME-Lite Karimireddy et al. (2020a) reinterpret federated learning as biased stochastic optimization and improve convergence by aligning federated updates with centralized SGD via momentum and gradient correction.
@@ 2.2 Parameter-Efficient Fine-Tuning and Federated LoRA
Parameter-efficient fine-tuning (PEFT) has become a standard approach for adapting large models under resource constraints. LoRA Hu et al. (2022) freezes the backbone and learns low-rank updates for selected layers, greatly reducing trainable parameters and communication costs, which makes it especially attractive in FL.
Several studies extend LoRA to federated settings. FedIT Zhang et al. (2024) performs FedAvg-style aggregation over LoRA parameters, enabling efficient collaboration but implicitly assumes homogeneous adapter configurations and can be sensitive to rank mismatch. To mitigate rank heterogeneity, recent work has explored stacking-based aggregation to align heterogeneous low-rank updates into a shared subspace. FLoRA Wang et al. (2024) is a representative approach that constructs an expanded low-rank subspace via stacking, providing a principled way to aggregate clients with different ranks without padding or truncation. Other variants improve initialization, aggregation bias, or fairness properties for federated LoRA training Bian et al. (2025).
These methods demonstrate that stacking is an effective tool for handling rank heterogeneity. However, they typically treat each client adapter as a monolithic update and implicitly assume that all low-rank directions should be globally aligned and aggregated. In federated LLM fine-tuning, client updates may contain a mixture of globally transferable knowledge and locally specific adaptations; indiscriminate alignment can over-regularize personalization and become fragile when additional constraints (e.g., privacy noise) are introduced.
@@ 2.3 Personalization and Privacy in Federated Learning
Personalized FL aims to account for client-specific objectives and data distributions by learning both shared and local components. Representative methods include Ditto Gao et al. (2024), which couples a global model with personalized local models via regularization, and FedRep Collins et al. (2021), which separates shared representations from client-specific heads. These approaches highlight the importance of explicitly distinguishing between globally shared and locally specialized parameters, but they are not designed around the structural properties of low-rank adapters in LLM fine-tuning.
@@ 2.4 Positioning of Our Work
In contrast to prior federated LoRA methods, our work focuses on the semantic roles of low-rank updates and the resulting need for structural awareness in aggregation. Rather than proposing another stacking mechanism, we introduce a dual-module LoRA decomposition that separates globally alignable updates from locally personalized ones. Stacking-based aggregation is then applied selectively to the global module, while local modules remain private and unaggregated. This perspective addresses not only how to aggregate heterogeneous ranks, but also which updates should be aggregated under personalization and privacy constraints, enabling robust federated LLM fine-tuning with improved stability under DP noise.
@@ 3.1 Preliminaries: Federated LoRA Fine-tuning
When ranks differ across clients, directly averaging low-rank factors is ill-defined without padding or truncation, and may introduce biased updates. Stacking-based aggregation alleviates dimensional mismatch by aligning heterogeneous updates into an expanded subspace, but implicitly assumes that all updates should be globally aligned and aggregated. In federated LLM fine-tuning, client updates often contain a mixture of globally transferable knowledge and client-specific adaptations; forcing global alignment on all updates can harm personalization and becomes fragile under differential privacy noise. Our goal is therefore to design an aggregation scheme that is robust to rank heterogeneity while respecting the semantic roles of low-rank updates.
3.2 Dual-Module LoRA Decomposition
@@
where $\Delta W_{k}^{(g)}$ is a global module intended to capture transferable directions shared across clients, and $\Delta W_{k}^{(l)}$ is a local module for client-specific adaptations.
This decomposition reflects an important observation in federated LLM fine-tuning: client-side low-rank updates are semantically heterogeneous. Some update directions correspond to transferable knowledge that should be aligned and shared across clients, while others encode client-specific or domain-specific adaptations whose forced alignment can be harmful. By explicitly separating $\Delta W_{k}^{(g)}$ and $\Delta W_{k}^{(l)}$, we impose a structural constraint on which subspaces are allowed to participate in cross-client aggregation.
Figure 1: Framework overview. Each client keeps a private (local) module and trains a public (global) adapter on a frozen LLM. Only the public module is communicated and aggregated via selective stacking; DP noise is applied to the public update.
As illustrated in Figure 1, we decouple transferable (public) updates from client-specific (private) adaptations, enabling selective aggregation and privacy-aware training.
@@
The two modules play different roles in the federated process: (i) the global module participates in cross-client aggregation and is broadcast each round; (ii) the local module is never aggregated and remains private to the client. This structural separation is the key design choice that allows us to restrict subspace alignment to semantically appropriate updates, rather than treating each LoRA adapter as a monolithic update.
@@
In each round $t$, client $k\in S_{t}$ receives the current global module $\theta^{(g,t)}$ (the collection of all $\{A^{(g)},B^{(g)}\}$ across adapted layers) and keeps its local module $\theta_{k}^{(l,t)}$. The client performs $E$ steps of local optimization to minimize:
@@
and uploads only the updated global module $\theta_{k}^{(g,t)}$ (or its update) to the server. The local module is kept on-device and continues across rounds.
3.3 Selective Stacking Aggregation on the Global Module
We propose selective stacking, a stacking-based aggregation strategy that is applied exclusively to the global module while explicitly excluding client-local updates from cross-client alignment. Let $\theta_{k}^{(g,t)}$ denote the global-module parameters trained by client $k$ at round $t$ for a given layer, i.e., $\theta_{k}^{(g,t)}=\{A_{k}^{(g,t)},B_{k}^{(g,t)}\}$ with rank $r_{k}^{(g)}$. To aggregate heterogeneous ranks without padding or truncation, we perform stacking-based aggregation only on the global module.
@@ 3.3 Selective Stacking Aggregation on the Global Module
$A^{(g)}_{t}=\mathrm{Concat}\!\left(\sqrt{p_{k}}\,A_{k}^{(g,t)}\right)_{k\in S_{t}}$ (4)
$B^{(g)}_{t}=\mathrm{Concat}\!\left(\sqrt{p_{k}}\,B_{k}^{(g,t)}\right)_{k\in S_{t}}$ (5)
@@
which is well-defined under rank heterogeneity. Importantly, local modules are excluded from stacking and aggregation, avoiding forced alignment of client-specific directions.
@@
Rank control via re-compression.
A practical consequence of stacking is that the effective rank of the aggregated global update grows with the number of participating clients and communication rounds. Concretely, the concatenation in Eqs. (5)–(6) expands the rank dimension, which increases both the transmission cost (larger adapter factors) and the memory footprint on the server and clients.
Moreover, under privacy-aware training, stochasticity (e.g., DP noise) and client heterogeneity tend to inject low-energy directions that accumulate across rounds, leading to unstable optimization and diminished generalization if left uncontrolled. To keep the global module within a fixed budget, we apply low-rank re-compression to $\Delta W_{t}^{(g)}$:
@@
In practice, we compute a truncated SVD (or an equivalent low-rank factorization) $\Delta W_{t}^{(g)}=U\Sigma V^{\top}$ and retain only the top-$r_{\max}$ singular components, yielding $\Delta W_{t}^{(g)}\approx U_{r}\Sigma_{r}V_{r}^{\top}$. We then convert this truncated form back to LoRA-style factors by setting $\tilde{B}_{t}^{(g)}=U_{r}\Sigma_{r}^{1/2}$ and $\tilde{A}_{t}^{(g)}=\Sigma_{r}^{1/2}V_{r}^{\top}$, so that $\tilde{B}_{t}^{(g)}\tilde{A}_{t}^{(g)}$ matches the best rank-$r_{\max}$ approximation under the Frobenius norm. This step acts as a spectral filter that preserves dominant, transferable directions while discarding tail components that are more likely to be noise-dominated, and it keeps the communicated global module compact for the next round. Finally, the server broadcasts $(\tilde{A}_{t}^{(g)},\tilde{B}_{t}^{(g)})$ as the next-round global module.
3.4 Privacy-Aware Optimization (DP Compatibility)
Our dual-module design is naturally compatible with differential privacy (DP). The key idea is to apply DP mechanisms only to the global module, while keeping local modules noise-free.
Client-side DP-SGD on the global module.
During local training, client $k$ clips per-sample (or per-microbatch) gradients for the global module and adds Gaussian noise:

### v2

# SDFLoRA: Selective Decoupled Federated LoRA for Privacy-preserving Fine-tuning with Heterogeneous Clients

@@ Abstract
Federated learning (FL) has emerged as a promising paradigm for adapting large language models (LLMs) to distributed data. To mitigate the high communication and memory overhead, parameter efficient techniques such as Low Rank Adaptation (LoRA) are widely adopted. However, practical deployments often exhibit rank and data heterogeneity, making direct aggregation of LoRA updates biased and unstable. Existing approaches enforce a unified rank or align heterogeneous updates into a single shared subspace, which undermines personalization and accuracy. Moreover, under differential privacy (DP), adding noise to such mixed updates could perturb client specific directions that should remain local, resulting in utility loss. To address these issues, we propose Selective Decoupled Federated LoRA (SDFLoRA), a structure aware LoRA framework that decouples each client update into a shared component for updating, and a private component that preserves client specific semantics. Subspace alignment and aggregation are applied selectively to the shared module, while private modules remain local. We further design low rank re-compression to maintain a fixed rank budget for the shared module. This structure supports privacy aware optimization by injecting DP noise exclusively into the shared module. Experiments on multiple benchmarks demonstrate that SDFLoRA beats federated LoRA baselines and exhibits strong robustness under privacy constraints.
@@ 1 Introduction
Parameter-efficient federated fine-tuning (PEFT) has emerged as a critical paradigm for adapting Large Language Models (LLMs) to distributed, privacy-sensitive datasets. By optimizing only lightweight adapters, clients avoid full-model communication and memory overhead, enabling deployments in domains such as healthcare, finance, and on-device intelligence. Among PEFT approaches, low-rank adaptation (LoRA) Hu et al. (2022) is the dominant choice in federated LLM adaptation and is widely adopted in recent FL-LLM systems Wang et al. (2024); Yang et al. (2025); long2024dual.
Despite these advances, practical Federated Learning (FL) deployments could exhibit strong heterogeneity, as clients originate from different domains, tasks, and user populations. While LoRA updates naturally contain both transferable directions shared across clients and client-specific directions that encode local semantics and personalization. However, existing federated LoRA pipelines typically enforce shared alignment and aggregation over all update directions, which suppresses client-specific adaptations and limits personalization. This issue is further exacerbated by rank heterogeneity, where clients adopt different low rank configurations due to diverse system constraints, making naïve aggregation even more biased and unstable.
Existing solutions for federated LoRA under heterogeneous ranks can be grouped into two paradigms. The first paradigm enforces a unified rank across all clients, which reduces flexibility and can waste resources on constrained devices Cho et al. (2024). The second paradigm aligns heterogeneous LoRA updates into a shared subspace to enable aggregation. However, it tends to align all directions indiscriminately, including client-specific components that should remain private and unshared, thereby harming personalization Zhao et al. (2023). Meanwhile, differential privacy (DP) is widely adopted in FL to protect private client information. When such structurally entangled updates are globally aligned, adding noise to the entire update could perturbs client specific components that are not meant to be shared, leading to utility loss and a privacy guarantee Wei et al. (2023).
Motivated by these observations, we propose Selective Decoupled Federated LoRA (SDFLoRA), which decomposes each client adapter into a shared module that captures transferable knowledge and a private module that preserves client specific adaptations. Only the shared module participates in selective subspace alignment and aggregation across clients, while private modules remain private and unaggregated. To control the rank growth induced by stacking-based aggregation, we further introduce low rank re-compression to maintain a fixed rank budget for the shared module. This structure naturally supports privacy aware optimization by applying DP mechanisms exclusively to the shared module updates, avoiding unnecessary perturbations on local adaptations.
@@
  • We propose SDFLoRA, a structure-aware federated LoRA framework that decouples each client adapter into a shared module and a private module, and performs selective subspace alignment and aggregation exclusively on the shared module. The shared module is further constrained by a fixed rank budget via low rank re-compression, enabling stable and efficient aggregation.
  • We introduce a DP compatible optimization enabled by the decoupling, where differential privacy noise is injected only into the aggregated shareable update. This design avoids unnecessary perturbations on client-specific directions and leads to improved robustness under privacy constraints.
  • Extensive experiments on multiple benchmarks with heterogeneous clients demonstrate that SDFLoRA outperforms representative federated LoRA baselines, particularly in achieving a superior balance between privacy protection and model performance.
@@ 2.1 Federated Optimization under Client Heterogeneity
FL enables collaborative model training without centralizing data by alternating between local optimization and server-side aggregation. Federated Averaging (FedAvg) McMahan et al. (2017) is the canonical approach, but its simple weighted averaging mechanism often leads to unstable convergence under non-IID data distributions, partial client participation, and multiple local update steps, a phenomenon commonly referred to as client drift.
To mitigate heterogeneity-induced optimization challenges, a lot of work has proposed algorithmic modifications from different perspectives. FedProx Li et al. (2020) introduces a proximal regularization term to constrain local updates from deviating excessively from the shared model. FedDyn Durmus et al. (2021) and FedDC Gao et al. (2022) further address client drift by incorporating dynamic correction terms that explicitly decouple local objectives from the shared optimization trajectory. Another line of work focuses on correcting biases caused by heterogeneous local computation. FedNova Wang et al. (2020) normalizes client updates according to their effective number of local steps, while SCAFFOLD Karimireddy et al. (2020b) reduces update variance through control variates shared between the server and clients. From an optimization-centric viewpoint, MIME and MIME-Lite Karimireddy et al. (2020a) reinterpret federated learning as biased stochastic optimization and improve convergence by aligning federated updates with centralized SGD via momentum and gradient correction.
@@ 2.2 Parameter-Efficient Fine-Tuning and Federated LoRA
PEFT has become a standard approach for adapting large models under resource constraints. LoRA Hu et al. (2022) freezes the backbone and learns low-rank updates for selected layers, greatly reducing trainable parameters and communication costs, which makes it especially attractive in FL.
Several studies extend LoRA to federated settings. FedIT Zhang et al. (2024) performs FedAvg-style aggregation over LoRA parameters, enabling efficient collaboration but implicitly assumes homogeneous adapter configurations and can be sensitive to rank mismatch. To mitigate rank heterogeneity, recent work has explored stacking-based aggregation to align heterogeneous low-rank updates into a shared subspace. FLoRA Wang et al. (2024) is a representative approach that constructs an expanded low-rank subspace via stacking, providing a principled way to aggregate clients with different ranks without padding or truncation. Other variants improve initialization, aggregation bias, or fairness properties for federated LoRA training Bian et al. (2025); guoselective.
These methods demonstrate that stacking is an effective tool for handling rank heterogeneity. However, they typically treat each client adapter as a monolithic update and implicitly assume that all low-rank directions should be sharedly aligned and aggregated. In federated LLM fine-tuning, client updates may contain a mixture of sharedly transferable knowledge and locally specific adaptations; indiscriminate alignment can over-regularize personalization and become fragile when additional constraints (e.g., privacy noise) are introduced.
@@ 2.3 Personalization and Privacy in Federated Learning
Personalized FL aims to account for client-specific objectives and data distributions by learning both shared and private components. Representative methods include Ditto Gao et al. (2024), which couples a shared model with personalized local models via regularization, and FedRep Collins et al. (2021), which separates shared representations from client-specific heads. These approaches highlight the importance of explicitly distinguishing between sharedly shared and locally specialized parameters, but they are not designed around the structural properties of low-rank adapters in LLM fine-tuning.
@@ 2.4 Positioning of Our Work
Compared to prior federated LoRA methods, our work focuses on the semantic roles of low-rank updates and the resulting need for structural awareness in aggregation. Rather than proposing another stacking mechanism, we introduce SDFLoRA that separates sharedly alignable updates from locally personalized ones. Stacking-based aggregation is then applied selectively to the shared module, while private modules remain private and unaggregated. This perspective addresses not only how to aggregate heterogeneous ranks, but also which updates should be aggregated under personalization and privacy constraints, enabling robust federated LLM fine-tuning with improved stability under DP noise.
@@ 3.1 Preliminaries: Federated LoRA Fine-tuning
When ranks differ across clients, directly averaging low-rank factors is ill-defined without padding or truncation, and may introduce biased updates. Stacking-based aggregation alleviates dimensional mismatch by aligning heterogeneous updates into an expanded subspace, but implicitly assumes that all updates should be sharedly aligned and aggregated. In federated LLM fine-tuning, client updates often contain a mixture of sharedly transferable knowledge and client-specific adaptations; forcing shared alignment on all updates can harm personalization and becomes fragile under differential privacy noise. Our goal is therefore to design an aggregation scheme that is robust to rank heterogeneity while respecting the semantic roles of low-rank updates.
3.2 Dual-Model LoRA Decomposition
@@
where $\Delta W_{k}^{(g)}$ is a shared module intended to capture transferable directions shared across clients, and $\Delta W_{k}^{(l)}$ is a private module for client-specific adaptations.
This decomposition reflects an important observation in federated LLM fine-tuning: client-side low-rank updates are semantically heterogeneous. Some update directions correspond to transferable knowledge that should be aligned and shared across clients, while others encode client-specific or domain-specific adaptations whose forced alignment can be harmful. By explicitly separating $\Delta W_{k}^{(g)}$ and $\Delta W_{k}^{(l)}$, we impose a structural constraint on the subspaces that are allowed to participate in cross-client aggregation.
Figure 1: Framework overview. Each client keeps a private module and trains a shared module on a frozen LLM. Only the shared module is communicated and aggregated via selective aggregation; Differential Privacy noise is applied to the public update.
As illustrated in Figure 1, we decouple transferable (public) updates from client-specific (private) adaptations, enabling selective aggregation and privacy-preserving training.
@@
The two modules play different roles in the federated process: (i) the shared module participates in cross-client aggregation and is broadcast each round; (ii) the private module is never aggregated and remains private to the client. This structural separation is the key design choice that allows us to restrict subspace alignment to semantically appropriate updates, rather than treating each LoRA adapter as a monolithic update.
@@
In each round $t$, client $k\in S_{t}$ receives the current shared module $\theta^{(g,t)}$ (the collection of all $\{A^{(g)},B^{(g)}\}$ across adapted layers) and keeps its private module $\theta_{k}^{(l,t)}$. The client performs $E$ steps of local optimization to minimize:
@@
and uploads only the updated shared module $\theta_{k}^{(g,t)}$ (or its update) to the server. The private module is kept on-device and continues across rounds.
3.3 Selective Stacking Aggregation on the Shared Module
We propose selective stacking, a stacking-based aggregation strategy that is applied exclusively to the shared module while explicitly excluding client-private updates from cross-client alignment. Let $\theta_{k}^{(g,t)}$ denote the shared-module parameters trained by client $k$ at round $t$ for a given layer, i.e., $\theta_{k}^{(g,t)}=\{A_{k}^{(g,t)},B_{k}^{(g,t)}\}$ with rank $r_{k}^{(g)}$. To aggregate heterogeneous ranks without padding or truncation, we perform stacking-based aggregation only on the shared module.
@@ 3.3 Selective Stacking Aggregation on the Shared Module
$A^{(g)}_{t}=\mathrm{Concat}\!\left(p_{k}\,A_{k}^{(g,t)}\right)_{k\in S_{t}}$ (4)
$B^{(g)}_{t}=\mathrm{Concat}\!\left(p_{k}\,B_{k}^{(g,t)}\right)_{k\in S_{t}}$ (5)
@@
which is well-defined under rank heterogeneity. Importantly, private modules are excluded from stacking and aggregation, avoiding forced alignment of client-specific directions.
@@
Rank control by re-compression.
A practical consequence of stacking is that the effective rank of the aggregated shared update grows with the number of participating clients and communication rounds. Concretely, the concatenation in Eqs. (5)–(6) expands the rank dimension, which increases both the transmission cost (larger adapter factors) and the memory footprint on the server and clients.
Moreover, under privacy-aware training, stochasticity (e.g., DP noise) and client heterogeneity tend to inject low-energy directions that accumulate across rounds, leading to unstable optimization and diminished generalization if left uncontrolled. To keep the shared module within a fixed budget, we apply low-rank re-compression to $\Delta W_{t}^{(g)}$:
@@
In practice, we compute a truncated SVD (or an equivalent low-rank factorization) $\Delta W_{t}^{(g)}=U\Sigma V^{\top}$ and retain only the top-$r_{\max}$ singular components, yielding $\Delta W_{t}^{(g)}\approx U_{r}\Sigma_{r}V_{r}^{\top}$. We then convert this truncated form back to LoRA-style factors by setting $\tilde{B}_{t}^{(g)}=U_{r}\Sigma_{r}^{1/2}$ and $\tilde{A}_{t}^{(g)}=\Sigma_{r}^{1/2}V_{r}^{\top}$, so that $\tilde{B}_{t}^{(g)}\tilde{A}_{t}^{(g)}$ matches the best rank-$r_{\max}$ approximation under the Frobenius norm. This step acts as a spectral filter that preserves dominant, transferable directions while discarding tail components that are more likely to be noise-dominated, and it keeps the communicated shared module compact for the next round. Finally, the server broadcasts $(\tilde{A}_{t}^{(g)},\tilde{B}_{t}^{(g)})$ as the next-round shared module.
3.4 Privacy Aware Optimization (DP Compatibility)
Our decoupled design is naturally compatible with differential privacy (DP). The key idea is to apply DP mechanisms only to the shared module, while keeping private modules noise-free.
Client-side DP-SGD on the shared module.
During local training, client $k$ clips per-sample (or per-microbatch) gradients for the shared module and adds Gaussian noise: